I need help please.
For a fundraising project, your Math Club is publishing an art calendar. The cost of the digital images and the permission to use it is P34, 000.00. In addition to these “one-time” charges, the unit cost of printing each calendar is P130.
a. Write a model that gives the average cost per calendar as a function of the number of calendars to be printed.
b. Graph the model manually or by using a graphing software or an online graphing tool and use the graph to estimate the number of calendars you need to print before the average cost drops to P200 per calendar.
c. Describe what happens to the average cost as the number of calendars printed increases.

Answer by Alex.33(110)

I have to ignore the 'P' in your terms like 'P34,000.00' and 'P130' because the meaning of them is unclear, and therefore have to assume you have omitted the units of money used.
a. the number of calendars and the average cost are represented by symbols x and p(x)
average cost per calendar =total cost/number of calendars
p(x)=(34000+130*x)/x=130+34000/x.
b. if you want to get a nice graph I recommand you take a look at https://www.desmos.com/:
And the result: .
The point on the function where p(x)=200 is already given. At this point the number of calendars printed is roughly 486.
c. it decreases. You are able to see it from the graph.